When mass is kept constant then terminal velocity for wood is greater or thr iron bob? Nd why
1
Expert's answer
2019-08-01T09:57:00-0400
"v_t\sim d(\rho_s-\rho)"

d is the diameter of the spherical object, ρs is the density of the object, ρ is the density of the fluid.


"d\sim (\frac{m}{\rho_s})^{\frac{1}{3}}"

Thus,


"v_t\sim \rho_s^{-\frac{2}{3}}"

Then terminal velocity for wood is greater than for iron bob.